### 14 Dec 2023 v1

# Aleth-NeRF: Illumination Adaptive NeRF with Concealing Field Assumption

Abstract
The standard Neural Radiance Fields (NeRF) paradigm employs a viewer-centered methodology, entangling the aspects of illumination and material reflectance into emission solely from 3D points. This simplified rendering approach presents challenges in accurately modeling images captured under adverse lighting conditions, such as low light or over-exposure. Motivated by the ancient Greek emission theory that posits visual perception as a result of rays emanating from the eyes, we slightly refine the conventional NeRF framework to train NeRF under challenging light conditions and generate normal-light condition novel views unsupervisedly. We introduce the concept of a ”Concealing Field,” which assigns transmittance values to the surrounding air to account for illumination effects. In dark scenarios, we assume that object emissions maintain a standard lighting level but are attenuated as they traverse the air during the rendering process. Concealing Field thus compel NeRF to learn reasonable density and colour estimations for objects even in dimly lit situations. Similarly, the Concealing Field can mitigate over-exposed emissions during rendering stage. Furthermore, we present a comprehensive multi-view dataset captured under challenging illumination conditions for evaluation. Our code and proposed dataset is available at https://github.com/cuiziteng/Aleth-NeRF.
Figure 1: Utilizing the Concealing Field assumption, Aleth-NeRF is capable of processing both low-light &\&& over-expose multi-view images as inputs and generating novel views with natural illumination.
Introduction
Can you see your days blighted by darkness ?
Pink Floyd “Lost of Words”
Neural Radiance Field (NeRF) (Mildenhall et al. 2020) has been demonstrated to effectively understand 3D scenes from 2D images and generate novel views. However, the formulation of NeRF and its follow-up variants assume captured images are under normal light, often failing to work under low-light (Mildenhall et al. 2021) or over-exposure scenarios. This is because vanilla NeRF is viewer-centered which models the amount of light emission from a location to the viewer without disentangling illumination and material (Fig. 1(a)) (Lyu et al. 2022). As a result, the NeRF algorithm interprets a dark scene as insufficient radiation of the 3D object particles, violating the estimation of the object’s material and geometry. In practical applications, images are often taken under challenging lighting conditions. Therefore, this paper aims to slightly modify vanilla NeRF for under &\&& over-exposure scenes. As shown in Fig. 1(c, d), the proposed, Aleth-NeRF, renders normal-light novel views despite the severe input images.
[…]
Our proposed method Aleth-NeRF takes low-light &\&& over-exposure images as inputs to train the model and learn the volumetric representation jointly with concealing fields. As shown in Fig.1(b), we jointly train NeRF with concealing fields between the object and viewer. For the low light scenario, we remove the concealing fields during the rendering stage (Fig. 1(c)). When dealing with over-exposure images, Aleth-NeRF would add concealing fields to suppress overly bright (Fig. 1(d)). Our contributions are summarized as follow:
  • We propose Aleth-NeRF, that trains under low-light &\&& over-exposure conditions and generates novel views under normal-light. Inspired by ancient Greek philosophy, we naturally extend the transmittance function in vanilla NeRF by modelling concealing fields between the objects and viewer to interpret lightness degradation.
  • We contribute a challenging illumination multi-view dataset, with paired sRGB low-light &\&& normal-light &\&& over-exposure images, dataset would also be public.
  • We compare with various image enhancement and exposure correction methods &\&& previous NeRF-based method (Mildenhall et al. 2021). Extensive experiments show that our Aleth-NeRF achieves satisfactory enhancement quality and multi-view consistency.
Related Works
Novel View synthesis with NeRF
NeRF (Mildenhall et al. 2020) is proposed for novel view synthesis from a collection of posed input images. The unique advantage of NeRF models exists in preserving the 3D geometry consistency thanks to its physical volume rendering scheme. In addition several methods have been proposed to speed up and improve NeRF training (Barron et al. 2021; Sara Fridovich-Keil and Alex Yu et al. 2022; Lindell, Martel, and Wetzstein 2021; Yu et al. 2021; Jain, Tancik, and Abbeel 2021; Deng et al. 2022; Müller et al. 2022).
Many of the latter works focus on improving NeRF’s performance under various degradation conditions, such as blurry (Ma et al. 2021), noisy (Pearl, Treibitz, and Korman 2022), reflection (Guo et al. 2022), glossy surfaces (Verbin et al. 2022), underwater (Levy et al. 2023), or use NeRF to handle super-resolution (Wang et al. 2021a; Bahat et al. 2022) and HDR reconstruction (Xin et al. 2021; Jun-Seong et al. 2022) in 3D space. Another line of research extends NeRF for lightness editing in 3D space. Some work, like NeRF-W (Martin-Brualla et al. 2021), focuses on rendering NeRF with uncontrolled in-the-wild images, other relighting works (Srinivasan et al. 2021; Rudnev et al. 2022; Zhang et al. 2021b) rely on known illumination conditions and introduce additional physical elements (i.e. normal, light, albedo, etc.), along with complex parametric modeling of these elements. Meanwhile, these methods are not specifically designed for low-light &\&& over-exposure conditions.
Among these, RAW-NeRF (Mildenhall et al. 2021) is more closer to our work, which proposes to render NeRF in HDR RAW domain and then post-process the rendered scene with image signal processor (ISP), RAW-NeRF has shown a preliminary ability to enhance the scene light but requires HDR RAW data for training, which make it hard to generalize on common used sRGB images. Instead our Aleth-NeRF could directly rendered on sRGB under &\&& over exposure images and injection unsupervised enhancement into 3D space by an effective concealing fields manner.
Enhancement in challenging light conditions
Challenging lightness can arise from multiple sources, encompassing natural lighting variances (such as low-light situations and overly bright scenes) as well as human-induced factors (like incorrect camera exposure settings). To tackle these challenge lighting conditions, numerous techniques for image enhancement and exposure correction have been developed and proposed.
Image Enhancement &\&& Exposure Correction: Image enhancement methods aims to enhance images with poor illumination, traditional methods usually rely on RetiNex theory (Land 1986; Guo, Li, and Ling 2017) or Histogram Equalization (Gonzalez and Woods 2006), currently deep neural networks (DNNs) based methods have become the mainstream solutions, series of CNN &\&& Transformer-based methods have been developed in this direction (Wei et al. 2018; Moran et al. 2020; Wang et al. 2021b, 2022; Jiang et al. 2021; Guo et al. 2020; Jin, Yang, and Tan 2022; Ma et al. 2022; Yang et al. 2023). Meanwhile, several exposure correction methods have been proposed to consider both under &\&& over exposure conditions (Afifi et al. 2021; Nsampi, Hu, and Wang 2021; Cui et al. 2022; Huang et al. 2023), which aims to correct underexposure and its adverse overexposure images into normal-light condition. However, image enhancement &\&& exposure correction methods almost build on 2D image space operations, which often fail to exploit the 3D geometry of the scene and could not deal with multi-view inputs.
Video Enhancement &\&& Burst Enhancement: Beyond above techniques that focus on single image. Video enhancement methods have been proposed to optimize the temporal consistency between adjacent frames, ensuring stability when processing different frames. These methods employ various approaches such as 3D convolution (Lv et al. 2018), optical flow (Zhang et al. 2021a), and event guidance (Liu et al. 2023). Burst enhancement also plays a crucial role in modern computational photography area (Delbracio et al. 2021; Hasinoff et al. 2016), where multiple frames are captured during exposure and processed using an ”align-merge-enhance” approach to produce a single output frame. In recent advancements, deep neural networks have been employed to replace traditional manual operation algorithms in these methods (Godard, Matzen, and Uyttendaele 2018; Mildenhall et al. 2018; Dudhane et al. 2022).
However, existing image &\&& video &\&& burst enhancement methods primarily focus on enhancing images in their original views, rather than generating coherent 3D scenes with novel views, for comparison we have to combine these enhancement methods with NeRF (see Table. 1 and Table. 2). In contrast, Aleth-NeRF is capable of directly synthesizing novel views under challenging light conditions while achieving state-of-the-art enhancement quality.
Methods
Figure 2: Train on adverse lighting condition images $\textbf{C}^{adv}$, Aleth-NeRF performs unsupervised lightness correction by (a). remove concealing fields in low-light conditions and (b). add concealing fields in over-exposure conditions.
Neural Radiance Field Revisited
Radiance Field is defined as the density $\sigma$ and RGB colour value $c$ of a 3D location x under a 2D viewing direction d. The density $\sigma$, on the one hand, represents the radiation capacity of the particle itself at x, and on the other hand, controls how much radiance is absorbed when other lights pass through x.
[…]
where
[…]
is known as the accumulated transmittance that denotes the radiance decay rate of the particle at $\textbf{r}(t)$ when it is occluded by particles closer to the camera (at $\textbf{r}(s)$, $s<t$). The integrals are computed by a discrete approximation over sampled 3D points along the ray $\bf r$ (see Fig. 1(b)). The discrete form of Eq. 1 and Eq. 2 are represented as follows:
${\bf C}({\bf r})=\sum_{i=1}^{N}T(\textbf{r}(i))(1-\exp(-\sigma(\textbf{r}(i))\cdot\delta))\cdot c(\textbf{r}(i),\textbf{d}),$ (3)
$T(\textbf{r}(i))=\exp\left(-\sum_{j=1}^{i-1}\sigma(\textbf{r}(j))\cdot\delta\right).$ (4)
Value of $\sigma(\textbf{r}(i))$ reflects the object occupancy in $\textbf{r}(i)$. $\delta$ is a constant distance value between adjacent sample points under uniform sampling.
Figure 3: Along camera ray r ($z$ axis), concealing fields and density $\sigma$ exhibit a negative correlation, $(x,y)$ denotes training images’ width and height.
Figure 4: A low-light “bike” scene for example, we show the ablation analyze of different loss functions’ effectiveness.
For the network structure, NeRF learns two multilayer perceptron (MLP) networks: density MLP $F_{\sigma}$ and colour MLP $F_{c}$, to map the 3D location $\textbf{r}(i)$ and 2D viewing direction d to its density $\sigma$ and colour $c$, specifically:
$F_{\sigma}(\textbf{r}(i))\rightarrow\sigma(\textbf{r}(i)),\textbf{h}$ (5)
$F_{c}(\textbf{h},\textbf{d})\rightarrow c(\textbf{r}(i),\textbf{d}),$ (6)
where h is a hidden feature vector that is sent to colour MLP $F_{c}$, color $c(\textbf{r}(i),\textbf{d})$ and density $\sigma(\textbf{r}(i))$ are further activated by Sigmoid and ReLU functions to regularise their ranges into $[0,1)$ and $[0,\infty)$ respectively. Given ground truth images C, NeRF is optimised by minimising the MSE loss between predicted images $\hat{\textbf{C}}$ and ground truth images C:
$\mathcal{L}_{mse}=\sum_{\textbf{r}}^{\textbf{R}}||\hat{\textbf{C}}(\textbf{r})-\textbf{C}(\textbf{r})||^{2}.$ (7)
We refer more details such as positional encoding and hierarchical volume sampling to NeRF paper (Mildenhall et al. 2020).
Aleth-NeRF with Concealing Field
Given adverse lighting condition images $\textbf{C}^{adv}(\textbf{r})$ taken in low-light &\&& over-exposure conditions, our goal is to generate novel views $\textbf{C}^{nor}(\textbf{r})$ in normal light condition. The key idea is that Aleth-NeRF assumes $\textbf{C}^{adv}$ and $\textbf{C}^{nor}$ are rendered with the same underlying density field $\sigma$ (yellow particle in Fig. 1) along each camera ray r, but with or without the proposed concealing fields (gray particle in Fig. 1).
We model two types of concealing fields to reduce light transport in the volume rendering stage: local Concealing Field denoted as $\Omega$ at the voxel level, and global Concealing Field denoted as $\Theta_{G}$ at the scene level. Multiplying two fields, $\Omega$ and $\Theta_{G}$ gives the final concealing field value which reduce the accumulated transmittance. Fig. 2 shows an overview of the Aleth-NeRF training strategy. Fig. 3 shows concealing fields $\Omega\cdot\Theta_{G}$’s distribution.
Local Concealing Field denoted as $\Omega(\textbf{r}(i))$, defines an extra light concealing capacity of a particle at 3D location $\textbf{r}(i)$. As depicted in Fig. 2, $\Omega$ is individually learned for each 3D position and is generated using the density MLP $F_{\sigma}$. To create the local concealing field $\Omega$, we introduce an additional large kernel convolution layer (with a size of 7) built upon $F_{\sigma}$. This convolution process establishes spatial relationships between pixels and enriches the Concealing Field with predominantly light-related information rather than structural information (Zamir et al. 2020). This larger kernel convolution also effectively suppresses noise and contributes to smoother rendering outcomes.
$conv_{size:7}(F_{\sigma}(\textbf{r}(i)))\rightarrow\Omega(\textbf{r}(i))$ (8)
Global Concealing Field denoted as $\Theta_{G}(i)$, is defined as a set of learnable parameters corresponding to the camera distance $i$ for all camera rays in R. $\Theta_{G}$ remains constant within a scene and is independent of voxels, as we posit that a specific degree of lighting influence remains consistent across the same scene. In our experiments, we initialize $\Theta_{G}(i)$ with a value of 0.5 for each camera distance $i$. With the global and local concealing fields, the accumulated transmittance $T$ in Eq. 4 is blighted by $\Omega$ and $\Theta_{G}$ to mimic the process of light suppression (see Fig. 1(b)), as follow:
$T^{conceal}(\textbf{r}(i))=\exp\left(-\sum_{j=1}^{i-1}\sigma(\textbf{r}(j))\cdot\delta\right)\cdot\prod_{j=1}^{i-1}\Omega(\textbf{r}(j))\Theta_{G}(j).$ (9)
Train and Test Schemes of Aleth-NeRF for both low-light and over-exposure scenes are illustrated in Fig. 2. For low-light conditions (blue arrow in Fig. 2), Aleth-NeRF incorporates $T^{conceal}$ during the volume rendering stage to generate low-light images $\hat{\textbf{C}}^{adv}$, while removing the concealing fields to produce normal-light images $\hat{\textbf{C}}^{nor}$ with the original $T$ in Eq. 4. Conversely, for over-exposure scenes (yellow arrow in Fig. 2), Aleth-NeRF employs $T$ to render very-bright images $\hat{\textbf{C}}^{adv}$ and incorporates the concealing fields along with $T^{conceal}$ to render normal-light images $\hat{\textbf{C}}^{nor}$. The NeRF regression loss is applied to compare predicted adverse-light images $\hat{\textbf{C}}^{adv}$ with the corresponding ground truth $\textbf{C}^{adv}$, while unsupervised lightness correction losses ensure the generation of predicted normal-light images $\hat{\textbf{C}}^{nor}$, more details please refer to next section.
Analysis of Concealing Fields is shown in Fig. 3, we analyse the distribution of concealing fields and density along the camera ray $\textbf{r}(i)$ ($z$ axis), where $x$, $y$ are the training images’ width and height. In Fig. 3, brighter region denotes the larger possibility there exists concealing fields (gray line) or density (red line). We find that concealing fields exist in location $\textbf{r}(i)$ with lower density value $\sigma(\textbf{r}(i))$. And the Pearson correlation coefficient Corr between concealing fields and density $\sigma$ are also negative along both $x-z$ axis (-0.589) and $y-z$ axis (-0.694). This validates that concealing fields are separated from density, thus rarely participating in scene rendering. Concealing fields exists more in locations $\textbf{r}(i)$ with sparse density, i.e. air outside the objects.
Effective Losses for Unsupervised Training
In this section, we present the loss functions that guide the unsupervised lightness correction of Aleth-NeRF, facilitating improve novel views synthesis and enhancement in low-light &\&& over-exposure conditions. An overview of the unsupervised training strategy is shown in Fig. 2.
NeRF Regression Loss: Direct apply NeRF’s MSE loss function $\mathcal{L}_{mse}$ in low-light and over-bright scenes would result in an imbalance of pixel weights. Dark pixels with small values would contribute minimally, while bright pixels with larger values would dominate the contribution. To rectify this imbalance and ensure effective NeRF training under non-standard lighting conditions, we propose the incorporation of an inverse tone curve (Brooks et al. 2019; Cui et al. 2021) denoted as $\Phi$. This curve serves to re-balance pixel weights, additionally we introduce a small value $\epsilon=1e^{-3}$ within the regression loss to facilitate more accurate estimations. As a result, the formulated inverse tone MSE loss $\mathcal{L}_{it-mse}$ for NeRF training is presented as follows:
$\Phi(x)=\frac{1}{2}-sin(\frac{sin^{-1}(1-2x)}{3}),$ (10)
$\mathcal{L}_{it-mse}=\sum_{\textbf{r}}^{\textbf{R}}||\Phi(\hat{\textbf{C}}(\textbf{r})+\epsilon)-\Phi(\textbf{C}(\textbf{r})+\epsilon)||^{2},$
the comparison of $\mathcal{L}_{mse}$ and $\mathcal{L}_{it-mse}$ is shown in Fig. 4(a), inverse tone MSE loss $\mathcal{L}_{it-mse}$ can enhance scene clarity while also suppress noise to a certain extent.
Degree &\&& Contrast Loss: In order to regulate the level of enhancement, we incorporate a degree loss denoted as $\mathcal{L}_{de}$ on the predict normal-light images $\hat{\textbf{C}}^{nor}$. $\mathcal{L}_{de}$ involves a hyperparameter $\mathbf{e}$, which signifies the enhancement degree inherent to Aleth-NeRF. To compute $\mathcal{L}_{de}$, a global average pooling operation is applied to the $\hat{\textbf{C}}^{nor}$ of each training batch. Subsequently, we minimize the degree loss $\mathcal{L}_{de}$ by comparing the pooled $\hat{\textbf{C}}^{nor}$ with the specified degree value $\mathbf{e}$. We set $\mathbf{e}$ to 0.4 in all the experiments, set $\mathbf{e}$ to other value in range of (0.35, 0.55) could also got different type desired enhancement results, ablation of $\mathbf{e}$ is in Fig. 4(b)
$\mathcal{L}_{de}=||\rm{avgpool}(\hat{\textbf{C}}^{nor}(\textbf{r}))-\mathbf{e}||^{2}.$ (11)
Then we additional design a contrast degree loss $\mathcal{L}_{co}$ between the adjacent sampling pixels of predict normal-light images $\hat{\textbf{C}}^{nor}$ and GT adverse-light images $\textbf{C}^{adv}$:
$\mathcal{L}_{co}=$ $\sum_{k\in[+1,-1]}(\hat{\textbf{C}}^{nor}(\textbf{r})-\hat{\textbf{C}}^{nor}(\textbf{r}+k))-$ (12)
$\mathbf{e}\cdot\eta\cdot(\textbf{C}^{adv}(\textbf{r})-\textbf{C}^{adv}(\textbf{r}+k)),$
the contrast degree loss would maintain the structure similarity between $\hat{\textbf{C}}^{nor}$ and $\textbf{C}^{adv}$, meanwhile control the contrast degree of $\hat{\textbf{C}}^{nor}$ by parameter $\eta$, ablation analyze of $\eta$ in low-light scene is shown in Fig. 4, when enhance degree $\mathbf{e}$ is fixed, higher $\eta$ represents higher contrast in $\hat{\textbf{C}}^{nor}$. While in over-exposure scene, $\mathbf{e}\cdot\eta$ would fixed to 2.
Color Constancy Loss: Photos taken in low-light &\&& over exposure conditions would easily lose their color information, directly remove &\&& add the concealing fields may cause color imbalance. To regularize the color of the predict normal-light images $\hat{\textbf{C}}^{nor}$, we introduce color constancy loss $\mathcal{L}_{cc}$ on $\hat{\textbf{C}}^{nor}$. Here we assume that $\hat{\textbf{C}}^{nor}$ obey the gray-world assumption (Buchsbaum 1980; Guo et al. 2020; Jin, Yang, and Tan 2022), as follows:
$\mathcal{L}_{cc}=\sum_{p,q}(\hat{\textbf{C}}^{nor}(\textbf{r})^{p}-\hat{\textbf{C}}^{nor}(\textbf{r})^{q})^{2},$ (13)
where $(p,q)\in\{(R,G),(G,B),(B,R)\}$ represents any pair of color channels, an ablation of color constancy loss $\mathcal{L}_{cc}$ can be found in Fig. 4(c).
Above all, Aleth-NeRF’s loss function $\mathcal{L}$ comprises four components: NeRF rendering loss $\mathcal{L}_{it-mse}$, along with three unsupervised lightness correction losses ($\mathcal{L}_{de}$, $\mathcal{L}_{co}$, and $\mathcal{L}_{cc}$). The overall training loss is then represented as:
$\mathcal{L}=\mathcal{L}_{it-mse}+\lambda_{1}\cdot\mathcal{L}_{de}+\lambda_{2}\cdot\mathcal{L}_{co}+\lambda_{3}\cdot\mathcal{L}_{cc},$ (14)
where $\lambda_{1}$, $\lambda_{2}$ and $\lambda_{3}$ are three non-negative parameters to balance total loss weights, which we set to $1e^{-3}$, $1e^{-3}$ and $1e^{-8}$ respectively.
Table 1: We assess novel view synthesis performance in low-light settings by comparing generated images with ground truth normal-light views. Our evaluation metrics include PSNR $\uparrow$ (indicating better quality), SSIM $\uparrow$ (indicating better structural similarity) and LPIPS $\downarrow$ (indicating closer perceptual similarity). Bold denotes the best result.
Method buu chair sofa bike shrub mean PSNR/ SSIM/ LPIPS PSNR/ SSIM/ LPIPS PSNR/ SSIM/ LPIPS PSNR/ SSIM/ LPIPS PSNR/ SSIM/ LPIPS PSNR/ SSIM/ LPIPS NeRF (Mildenhall et al. 2020)  7.53/ 0.313/ 0.414  6.01/ 0.142/ 0.599 6.26/ 0.206/ 0.566  6.32/ 0.067/ 0.625  8.00/ 0.027/ 0.684  5.24/ 0.151/ 0.578 ① NeRF + Image Enhancement Methods NeRF + RetiNexNet (Wei et al. 2018) 16.18/ 0.754/ 0.385 16.74/ 0.731/ 0.531 15.90/ 0.826/ 0.512 17.79/ 0.654/ 0.552 15.24/ 0.283/ 0.577 16.37/ 0.657/ 0.495 NeRF + Zero-DCE (Guo et al. 2020) 14.90/ 0.678/ 0.578 15.11/ 0.615/ 0.658 17.28/ 0.713/ 0.624 14.26/ 0.547/ 0.582 12.40/ 0.384/ 0.647 14.79/ 0.587/ 0.618 NeRF + SCI (Ma et al. 2022) 13.66/ 0.769/ 0.433 14.42/ 0.736/ 0.545 18.82/ 0.813/ 0.569 12.61/ 0.560/ 0.550 16.87/ 0.411/ 0.555 15.27/ 0.658/ 0.530 NeRF + IAT (Cui et al. 2022) 14.00/ 0.652/ 0.421 14.03/ 0.706/ 0.562 10.58/ 0.532/ 0.668 16.30/ 0.660/ 0.532 10.03/ 0.149/ 0.577 12.99/ 0.540/ 0.552 ② Image Enhancement Methods + NeRF RetiNexNet (Wei et al. 2018) + NeRF 16.16/ 0.777/ 0.339 16.82/ 0.773/ 0.488 16.87/ 0.806/ 0.548 18.00/ 0.717/ 0.448 14.65/ 0.269/ 0.513 16.50/ 0.668/ 0.467 Zero-DCE (Guo et al. 2020) + NeRF 13.81/ 0.783/ 0.353 11.20/ 0.694/ 0.498 12.02/ 0.747/ 0.420 11.35/ 0.586/ 0.448 14.11/ 0.426/ 0.473 12.50/ 0.647/ 0.438 SCI (Ma et al. 2022) + NeRF 13.62/ 0.821/ 0.309 11.75/ 0.756/ 0.526 10.10/ 0.750/ 0.505 18.10/ 0.644/ 0.458 17.13/ 0.500/ 0.469 14.14/ 0.694/ 0.453 IAT (Cui et al. 2022) + NeRF 14.33/ 0.697/ 0.311 14.77/ 0.704/ 0.563 17.77/ 0.811/ 0.519 13.68/ 0.621/ 0.501 13.84/ 0.322/ 0.536 14.88/ 0.631/ 0.486 ③ Video Enhancement Methods + NeRF MBLLEN (Lv et al. 2018) + NeRF 17.34/ 0.617/ 0.493 16.24/ 0.698/ 0.502 13.69/ 0.611/ 0.488 17.41/ 0.512/ 0.550 13.17/ 0.501/ 0.555 15.57/ 0.588/ 0.518 LLVE (Zhang et al. 2021a) + NeRF 15.54/ 0.755/ 0.473 15.46/ 0.756/ 0.492 17.89/ 0.801/ 0.457 15.47/ 0.699/ 0.455 14.97/ 0.414/ 0.511 15.87/ 0.685/ 0.470 ④ Our Proposed End-to-end Method Aleth-NeRF 19.14/ 0.839/ 0.306 16.96/ 0.793/ 0.483 16.97/ 0.847/ 0.367 17.56/ 0.719/ 0.468 17.55/ 0.484/ 0.451 17.64/ 0.736/ 0.415
Experiments
In this section, we first introduce our collected dataset with various exposure paired multi-view images, then we present the novel view synthesis results of Aleth-NeRF under both low-light and under-exposure settings. For practical implementation, our framework builds on the open-source PyTorch toolbox NeRF-Factory  [^2] https://github.com/kakaobrain/nerf-factory. We utilize the Adam optimizer with an initial learning rate of $5e^{-4}$ and employ a cosine learning rate decay strategy every 2500 iterations. The training batch size is set at 4096 for a total of 62500 iterations. To ensure fairness, all comparative experiments within the proposed dataset use the same training strategies and parameter configurations.
Challenging Illumination Multi-view Dataset
In this section, we introduce our collected paired low-light &\&& normal-light &\&& over-exposure multi-view dataset. Previous work (Mildenhall et al. 2021) also includes some low-light scenes, but their dataset concentrates on RAW denoising rather than sRGB low-light enhancement and does not include normal-light sRGB ground truth, making it hard to evaluate enhancement performance in novel view synthesis.
We collect 5 scenes (“buu”, “chair”, “sofa”, “bike”, “shrub”) in real-world. Each scene includes $25\sim 65$ images, for each image, we generate low-light &\&& normal-light &\&& over-exposure pairs by adjusting exposure time and ISO while other configurations of the camera are fixed. We capture multi-view images by moving and rotating the camera tripod. Images are collected with resolution 3000 $\times$ 4000. We down-sample the original resolution with ratio 8 to 375 $\times$ 500 for convenience, and generate the ground truth view and angle information by adopting COLMAP (Schönberger and Frahm 2016; Schönberger et al. 2016) on the normal-light part images. For dataset split, in each scene, we choose 3 $\sim$ 5 images as the testing set, 1 image as the validation set, and other images to be the training set. More details of our dataset please refer to supplementary.
Figure 5: An comparison of enhancement results and model efficiency with RAW-NeRF (Mildenhall et al. 2021), note that RAW-NeRF take 16-bits HDR image as inputs while NeRF &\&& Aleth-NeRF take 8-bits LDR image as inputs.
Figure 6: Novel view synthesis comparison in low-light and over-exposure conditions.
Novel View Synthesis in Low-light Condition
In this section, we show novel view synthesis results in low-light scenes, and make comparison between predicted normal-light views with ground truth normal-light views (see Table. 1). We design multiple comparison experiments in proposed dataset as follow:
We start by training vanilla NeRF (Mildenhall et al. 2020) under low-light conditions as a baseline (first row in Table 1). ① Subsequently, we apply various state-of-the-art enhancement methods to post-process the rendered NeRF results (“NeRF + ” in Table 1). ② Then we pre-process low-light images with enhancement methods and then train NeRF on these enhanced images (“* + NeRF” in Table. 1). Notably, we observe that the “NeRF + *” methods tend to underperform compared to the “* + NeRF” methods, especially in perceptual similarity (LPIPS $\downarrow$). This potentially arise from NeRF’s production of poor-quality images in dark scenes, causing enhancement methods to struggle when applied to the generated dark scene by NeRF. ③ At last, we apply two video enhancement methods MBLLEN (Lv et al. 2018) and LLVE (Zhang et al. 2021a) to pre-process low-light images and then trained with NeRF. ④ Above all, our Aleth-NeRF could gain the best performance on most scenes (see Table. 1), and trained in an end-to-end manner without any pre-processing nor post-processing.
Beyond enhancement methods, RAW-NeRF (Mildenhall et al. 2021) also conducted experiments under low-light conditions, due to their model is trained on RAW data and hard to implement on our dataset, a comparison with RAW-NeRF on their proposed dataset scene is shown in Fig. 5, while NeRF and Aleth-NeRF are trained with 8-bits LDR inputs and RAW-NeRF is trained with 16-bits HDR inputs, our Aleth-NeRF could also gain delight enhancement results while save much GPU memory and training time.
Fig. 6 shows qualitative visualization results in “sofa” and “shrub” scenes, with the comparison of current SOTA image enhancement methods Zero-DCE (Guo et al. 2020), SCI (Ma et al. 2022) and video enhancement method LLVE (Zhang et al. 2021a), we can see that our Aleth-NeRF could gain more vivid enhancement results and closer to ground truth, meanwhile also gain much better in multi-view consistency (see depth maps). More visualization results please refer to supplenmentary.
Table 2: Novel view synthesis performance in over-exposure setting by comparing generated images with ground truth normal-light views. We show the mean value of all 5 scenes.
Method PSNR/ SSIM/ LPIPS Method PSNR/SSIM/LPIPS NeRF + HE 12.72/ 0.553/ 0.588 HE + NeRF 15.45/ 0.647/ 0.558 NeRF + MSEC 12.98/ 0.514/ 0.605 MSEC + NeRF 13.89/ 0.558/ 0.541 NeRF + IAT 11.77/ 0.541/ 0.558 IAT + NeRF 14.31/ 0.536/ 0.507 NeRF 6.74/ 0.285/ 0.867 Aleth-NeRF 15.68/ 0.704/ 0.512
Novel View Synthesis in Over-Exposure Condition
Then we show novel view synthesis results in over-exposure scenes, similar to low-light, we choose 3 exposure correction methods to make comparison, including histogram equalization (HE) (Gonzalez and Woods 2006) and two SOTA deep network methods MSEC (Afifi et al. 2021) and IAT (Cui et al. 2022). As it shown in Table. 2, we use vanilla NeRF trained on over-exposure images to serve as a baseline, and adopt exposure correction methods on generated results (denote as “NeRF + *”). Also adopt exposure correction methods to pre-process over-exposure images and then train NeRF on the processed images (denote as “* + NeRF”). Overall, Aleth-NeRF could gain best results in PSNR and SSIM and second best result in LPIPS. Some visualization results are shown in Fig. 6, our Aleth-NeRF could suppress over-bright meanwhile maintain multi-view consistency.
Conclusion
In this paper, we propose Aleth-NeRF model. Inspired by wisdom of ancient Greeks, we introduce a concept: Concealing Field. After incorporating the concealing fields into NeRF, we can effectively address novel view synthesis tasks in scenes with low-light &\&& over-exposure condition.
One limitation is that Aleth-NeRF should be specifically trained for each scene, which is the same as vanilla NeRF. Besides, Aleth-NeRF may fail in scenes with non-uniform lighting conditions (Wang, Xu, and Lau 2022) or shadow conditions (Qu et al. 2017), we’ll solve these in the future.
Acknowledgement
This work was done during first author internship at Shanghai Artificial Intelligence Laboratory, means the corresponding author. This work was partially supported by JST Moonshot R&\&&D Grant Number JPMJPS2011, CREST Grant Number JPMJCR2015 and Basic Research Grant (Super AI) of Institute for AI and Beyond of the University of Tokyo. And this work is supported in part by the National Key R&\&&D Program of China (NO. 2022ZD0160100).